1/4 x² - 1/4 xy + 1/16 y² Factor it


1
Expert's answer
2021-10-21T14:43:25-0400

Find the greatest common factor of the terms

"\frac{1}{4}"


let's focus on:

"\frac{1}{4}x^2-\frac{1}{4}xy+\frac{1}{16}y^2"


Separate out "\frac{1}{4}"​ from each term

"\frac{1}{4}(x^2)+ \frac{1}{4}(-yx)+ \frac{1}{4}(\frac{y^2}{4})"


Factor out the greatest common factor

"\frac{1}{4}(x^2-yx+\frac{y^2}{4})"


Factor

"\frac{1}{4}(x-\frac{y}{2})^2"